# Random Forests as Statistical Procedures: Design, Variance, and Dependence

### v2

@@ 6.5 Split-candidate restriction size $q$
Lemma 1 does not assert that trees select the same splits, but isolates the design mechanism by which candidate restriction controls the opportunity for repeated split selection across trees. Whether this opportunity translates into actual alignment depends on the stability of split selection, which is addressed by the following result.
@@
where $Z_{\theta}^{(q)}(x)=\mathbb{E}[T_{\theta}^{(q)}(x)\mid\mathcal{I}_{\theta},X]$ denotes the resampling-conditional prediction rule under candidate-set size $q$, and $\theta,\theta^{\prime}$ are independent draws from the tree-generating mechanism. Assume that at each internal node along the path to $x$, the splitting criterion has a unique maximizer over the candidate set almost surely. Then
@@
Proposition 2 formalizes the role of candidate restriction as a mechanism for weakening alignment-induced dependence. Reducing $q$ decreases the probability that independently generated trees induce identical local averaging rules at $x$, thereby reducing the alignment component of the structural covariance.
Candidate restriction also affects the induced predictor itself through its impact on the distribution of local averaging rules. Because the infinite-aggregation predictor $f_{\infty}(x)$ is fully determined by the distribution of the resampling-conditional prediction rule $Z_{\theta}(x)$, any change in candidate-set size that alters this distribution necessarily alters the induced predictor.
Remark 6 (Effect of candidate restriction on the induced predictor).
The infinite-aggregation predictor $f_{\infty}(x)$ is the expectation of the prediction rule $Z_{\theta}(x)$. If two designs induce different distributions of $Z_{\theta}(x)$, their expectations differ unless $Z_{\theta}(x)$ is almost surely constant, which occurs only in degenerate cases. Consequently, changes in $q$ that alter the distribution of $Z_{\theta}(x)$ necessarily alter the induced predictor.
Resolution at $x$ is governed by the average weight vector $\mathbb{E}[W(x;\theta)\mid X]$, which depends on the distribution of the resampling-conditional prediction rule $Z_{\theta}(x)$. Increasing $q$ enlarges the shared candidate information across trees, concentrating the distribution of induced local averaging rules and reducing variability in split selection along the path to $x$. Conversely, decreasing $q$ increases heterogeneity in the induced neighborhoods and broadens the distribution of weight vectors. Thus, larger $q$ produces a more stable and concentrated resolution structure, while smaller $q$ increases diversity of local averaging behavior.
In contrast, reducing $q$ weakens alignment and decreases alignment-induced dependence, while simultaneously increasing randomness in the induced local averaging rules through greater variability in split selection. Candidate restriction redistributes structure between dependence and diversity of local averaging behavior, rather than eliminating either source of structure, and is central to the practical success of random forests.
6.6 Minimum terminal occupancy $s$
The minimum terminal occupancy parameter $s$ governs the spatial resolution of the tree-level prediction rules by imposing a lower bound on the number of observations contained in each terminal region. Larger values of $s$ enforce coarser partitions by preventing further splitting once terminal regions contain fewer than $s$ observations, while smaller values permit continued splitting and therefore finer localization around the prediction point $x$. Because the infinite-aggregation predictor $f_{\infty}(x)$ is determined by the distribution of terminal regions (equivalently, by the induced distribution of weight vectors), changes in $s$ act by modifying the scale at which local averaging is performed.
Increasing $s$ enlarges the terminal regions used to form predictions, producing coarser local averaging behavior and reduced sensitivity to fine-scale variation in the observed outcomes. Decreasing $s$ permits smaller terminal regions and therefore higher-resolution predictors that adapt more sharply to local features of the realized data. This effect is structural: even under infinite aggregation, the induced predictor reflects the resolution attainable by trees grown subject to the terminal-occupancy constraint.
The same mechanism also affects variability and dependence across trees. Smaller terminal regions average fewer observations, which increases sensitivity to local outcome variation and raises the marginal dispersion of individual tree predictions. Larger values of $s$ stabilize individual trees by enforcing broader averaging, but simultaneously promote reuse of similar coarse partitions across tree realizations. This increased regularity strengthens alignment of induced local averaging rules and can therefore increase structural dependence across trees.
Proposition 3 (Terminal occupancy and local resolution).
Consider two forest designs that differ only in the minimum terminal occupancy, with $s^{\prime}<s$. Under the monotone-growth assumption, reducing $s$ shifts the distribution of terminal regions at $x$ toward smaller local averaging sets.
Proposition 3 formalizes the role of the terminal-occupancy constraint as a resolution parameter. Lowering $s$ allows finer partitioning near $x$, while increasing $s$ enforces coarser averaging that persists under infinite aggregation. The proof relies only on the structural nesting of partitions induced by monotone tree growth and does not require distributional assumptions on the outcomes.
@@ 6.7 Sample size and persistence of structural dependence
This subsection addresses whether increasing sample size alone can eliminate structural dependence across trees when such dependence is induced by the tree-generating design. Throughout, the design parameters $(p_{\mathrm{obs}},q,s)$ are held fixed as $n$ varies.
Write
$f_{\infty,n}(x)=\mathbb{E}[T_{\theta}^{(n)}(x)\mid X],\qquad C_{T,n}(x)=\operatorname{Cov}\!\left(T_{\theta}^{(n)}(x),T_{\theta^{\prime}}^{(n)}(x)\mid X\right),$
where $T_{\theta}^{(n)}$ denotes a tree grown on a dataset of size $n$ under the same randomized construction.
Existing large-sample analyses of random forests describe how increasing $n$ refines the induced partitions and improves the resolution of the infinite-aggregation predictor. These results concern how the local averaging behavior at $x$ evolves with additional data, but they do not address whether dependence between independently generated trees weakens as the sample size grows.
From the finite-sample covariance decomposition, $C_{T,n}(x)$ is governed by the joint distribution of resampling patterns and induced local averaging rules. Increasing $n$ refines these rules but does not alter the sources of randomization in the tree-generating mechanism. In particular, when split selection along the path to $x$ remains randomized—due to candidate restriction, enforced terminal occupancy, or non-unique split choices—the distribution of local prediction rules does not collapse to a single deterministic rule.
Consequently, increasing sample size alone does not remove design-induced dependence across trees. Any reduction in structural covariance must arise from changes to the design itself, not from refinement of partitions driven by larger $n$.
@@ 6.8 Why added randomization improves finite-sample behavior
Additional randomization in split selection is a design parameter that directly alters the dependence structure of the forest and the distribution of induced local prediction rules. Unlike aggregation, which averages independent realizations of a fixed randomized procedure, added randomization modifies the tree-generating mechanism itself and therefore changes both the induced predictor and its structural dependence properties.
@@
A concrete example is augmented bagging, which introduces $r\geq 0$ independent noise variables into the predictor space prior to growing each tree, making $r$ an explicit randomization parameter (Mentch and Zhou, 2022). Increasing $r$ reduces the probability that any particular split is repeatedly selected along the path to a prediction point $x$, thereby dispersing the distribution of local decision rules induced by the forest design.
From the perspective developed in Section 4.3, this dispersion weakens alignment of induced prediction rules across trees. When randomization is weak, the tree-generating mechanism can concentrate probability on a small family of local averaging rules at $x$, leading independently generated trees to repeatedly induce similar neighborhoods and hence strong alignment-induced dependence. Increasing the level of randomization reduces this concentration by lowering the chance that the same early split decisions recur across trees, thereby weakening alignment and reducing the corresponding contribution to structural dependence.
At the same time, added randomization necessarily alters the distribution of induced partitions and local averaging behavior. By spreading probability mass across a broader collection of admissible decision rules, increased randomization modifies the resolution of the induced predictor at $x$. Local averaging behavior becomes more diverse across tree realizations, reflecting greater variability in how observations are grouped with $x$ under the design.
These two effects act in opposition. Added randomization weakens alignment and reduces structural dependence across trees, while simultaneously redistributing probability mass across local prediction rules and altering resolution. Improvements in finite-sample behavior arise when the reduction in alignment-induced dependence outweighs the accompanying change in resolution. This trade-off is inherently finite-sample: increasing the number of trees suppresses Monte Carlo variability but does not modify the design-induced structure that governs alignment and resolution.

### v3

@@ 6.3 Split-candidate restriction size $\boldsymbol{\mathrm{(}q)}$
A proof is provided in Appendix A.3. Lemma 1 isolates the design mechanism by which candidate restriction controls the opportunity for repeated split selection across trees. The following result shows that this increased opportunity translates directly into stronger alignment.
@@
where $Z_{\theta}^{(q)}(x)=\mathbb{E}[T_{\theta}^{(q)}(x)\mid\mathcal{I}_{\theta},X]$ denotes the resampling-conditional predictor under candidate-set size $q$, and $\theta,\theta^{\prime}$ are independent draws from the tree-generating mechanism. Assume that at each internal node along the path to $x$, the splitting criterion selects a single best split almost surely. Then
@@
A proof is provided in Appendix  A.4. Reducing $q$ therefore decreases the probability that independently generated trees lead to similar conditional predictors at $x$, reducing the alignment component of the $C_{T}(x)$.
Candidate restriction simultaneously affects resolution. Increasing $q$ concentrates the distribution of induced conditional predictors by reducing variability in split selection, producing more stable resolution at $x$. Decreasing $q$ increases diversity across tree realizations but broadens the weight distribution. Unlike $p_{\mathrm{obs}}$, which acts exclusively through joint inclusion probabilities, candidate restriction also affects the observation-overlap component indirectly: when trees make similar splitting decisions, a jointly included observation is more likely to receive positive weight at $x$ in both trees. Candidate restriction therefore modulates both terms of the covariance decomposition (3).
Figure 4: Quantile profiles of the true covariance floor $C_{T}(x)$ across $n_{\mathrm{test}}=400$ test points (Scenario: $n=400$, $p=10$; DGM in Appendix B). Left column varies the observation sampling scheme at fixed $q=4$; right column varies $q$ at fixed bootstrap resampling. Top row: continuous outcomes; bottom row: binary outcomes. Increasing $q$ increases $C_{T}(x)$. Increasing $p_{obs}$ increases $C_{T}(x)$, consistent with the increased tree correlation induced by larger candidate sets.
6.4 Minimum terminal occupancy $\boldsymbol{\mathrm{(}s)}$ and maximum tree depth
The minimum terminal node occupancy $s$ imposes a lower bound on the number of observations in each terminal region, defining the scale at which local averaging may be realized. Larger values of $s$ lead to coarser partitions with greater stability but reduced resolution; smaller values lead to finer partitioning and higher-resolution predictors at the cost of increased single-tree variability. Because coarser partitions are more likely to be reused across tree realizations, increasing $s$ strengthens alignment-induced dependence and raises $C_{T}(x)$ relative to a forest with smaller terminal nodes but the same candidate set size $q$. This effect is structural and persists under infinite aggregation. A maximum tree depth constraint $d_{\max}$ operates through the same mechanism: shallower trees produce coarser partitions, increasing between-tree agreement and inflating $C_{T}(x)$. When both constraints are active, the effective resolution at xx—and therefore the magnitude of the covariance floor—is determined by whichever is more restrictive.
@@ 6.5 Sample size and persistence of structural dependence
We now address how increasing sample size affects the covariance floor when the design parameters $(p_{\mathrm{obs}},q,s)$ are held fixed.
From the finite-sample covariance decomposition, $C_{T,n}(x)$ is governed by the joint distribution of resampling patterns and induced conditional predictors. Increasing $n$ has two competing effects. First, larger samples improve the quality of each tree’s local estimate: terminal nodes contain more observations, reducing single-tree variance and shrinking the Monte Carlo component $B^{-1}\sigma^{2}_{T,\theta}(x)$. Second, with more data each tree estimates $\mu(x)$ more accurately, reducing the between-tree dispersion that drives $C_{T}(x)$. This tighter agreement across trees can reduce $C_{T}(x)$ in absolute terms, as will be shown empirically in Section 8 where $\overline{C_{T}}$ is consistently smaller at $n=400$ than at $n=200$ for matched $(p,q,p_{\mathrm{obs}})$ configurations. However, increasing $n$ does not eliminate structural dependence. The covariance floor decreases but the design-induced randomization that produces it—candidate restriction, enforced terminal occupancy, non-unique split choices—is unchanged, so $C_{T}(x)$ remains strictly positive for any finite $n$.
@@ 6.6 Why added randomization improves finite-sample behavior
Additional randomization in split selection directly alters the dependence structure of the forest. Unlike aggregation, which averages independent realizations of a fixed randomized procedure, added randomization modifies the tree-generating mechanism itself and therefore changes both the induced predictor and its structural dependence properties.
@@
A concrete example is augmented bagging, which introduces $r\geq 0$ independent noise variables into the predictor space prior to growing each tree (Mentch and Zhou, 2022). Increasing $r$ reduces the probability that any particular split is repeatedly selected along the path to $x$, dispersing the distribution of conditional predictors induced by the forest design. From the perspective of Section 4.3, this weakens alignment across trees. At the same time, added randomization spreads probability mass across a broader collection of conditional prediction rules, altering the resolution of the induced predictor.